OpenStudy (anonymous):

For this Equation: √x-4 + 3 =9 My Answer Was x=40, Am I Correct?

OpenStudy (anonymous):

it'd be less irritating if you stop putting capital letters at the front of each word

OpenStudy (anonymous):

is that \[\sqrt{x-4}\] or \[\sqrt{x} - 4\]

OpenStudy (anonymous):

Lol Its Kindof A Habit...But Its The First One

OpenStudy (anonymous):

you got it and check is easy \[\sqrt{40-6}+3=\sqrt{36}+3=6+3\]

OpenStudy (anonymous):

oh ok, im just finding it really annoying sorry?

OpenStudy (anonymous):

You know it's only meant to be on the first word of every sentence right? Not Every Word In Every Sentence You Write Ever. see how bad that looks?

OpenStudy (anonymous):

you only do it if you're writing a title of something.

OpenStudy (anonymous):

okay sorry is that better,, and yes I know that...

OpenStudy (anonymous):

much better, im even smiling... i hope i didn't offend you doing that so extensively

OpenStudy (anonymous):

umm...no its okay.thats just something that I do ...I never really thought anything of it.

OpenStudy (anonymous):

oh ok, well other people might say worse things about it than just 'stop in a really long way'. :)

OpenStudy (anonymous):

yeah i understand